# Formation of Representations in Neural Networks

### v1

@@ 7.2 Feature-Importance Alignment and Invariant Learning
Another direct interpretation of $\mathbb{E}[gg^{\top}]$ is the saliency of a latent feature. The quantity $\mathbb{E}[gg^{\top}]$ measures how the loss function changes as a neuron activation is changed by a small degree and is often used on the input layer to measure the importance of a feature (Selvaraju et al.,, 2017; Adadi and Berrada,, 2018). This interpretation justifies many heuristics for understanding neural networks: take the principal components of a hidden layer and study its largest eigenvalues. If there is no direct link between the magnitude of the eigenvalues in these latent representations and their importance in affecting the learning (measured by the training loss), these analyses will be unreasonable.
@@
Corollary 1.
When CRH holds for a layer with activations $h_{a}$ and $h_{b}$, $f(x)$ is invariant to infinitesimal transforms of $x$ if and only if $h_{b}(x)$ is invariant to infinitesimal transforms of $x$.
In some sense, Proposition 1 can be seen as a generalization of NC (next section) because it applies to both regression and classification tasks. Note that there are two ways for the condition $\ell(f(h+\epsilon\hat{n}))=\ell(f(h))$ to be satisfied: (1) $f(h+\epsilon\hat{n})=f(h)$, which means that when the model output itself invariant to such a variation, the latent representation in this direction will vanish; (2) $\ell(f(h)+\epsilon\hat{n}^{T}\nabla f(h))=\ell(f(h))$, which means that any variation that does not change the loss function value will vanish. This fact also means that the model will learn a compact representation: any irrelevant latent space will have zero variation. This is consistent with the often observed matching between the rank of the representation and the inherent dimension of the problem (Papyan,, 2018; Ziyin,, 2024). A major phenomenon of well-trained CNNs is that the latent representations learn to become essentially invariant to task-irrelevant features (Zeiler and Fergus,, 2014; Selvaraju et al.,, 2017), an observation that matches the high-level features of the human visual cortex (Booth and Rolls,, 1998). The CRH thus suggests a reason for these observed invariances.
@@ 7.3 Reduction to Neural Collapse in Perfect Classification
We now prove that NC is equivalent to the CRH in an interpolating classifier. We focus on the first two properties here because NC3-4 are essentially consequences of NC1-2 (Rangamani and Banburski-Fahey,, 2022). NC1 states that the inner-class variations of the penultimate representation vanishes. NC2 states that the average representation $\mu_{c}$ of the class $c$ is orthogonal to each other: $\mu_{c}^{\top}\mu_{c^{\prime}}=\delta_{cc^{\prime}}$. The ground truth model must be invariant to inner class variations in a classification task. Let $c$ denote the index of a class, it should be the case that the ground truth model $f^{\prime}$ satisfies: $f^{\prime}_{c}(x_{c})=\zeta\mathbf{1}_{c}$, where $\mathbf{1}_{c}$ is a one hot vector on the $c$-th dimension and $\zeta>0$ is an arbitrary scalar. Proposition 1 thus suggests that any model that recovers the ground truth must have such invariance in the output and in the latent representation $h$, which is exactly NC1.
@@ 7.4 Equivariance of CRH and Neural Feature Ansatz
A closely related hypothesis is the NFA, which is related to the feature learning in fully connected layers (Radhakrishnan et al.,, 2023; Beaglehole et al.,, 2023). Using our notation, the NFA states $W^{\top}W\propto\mathbb{E}\left[\nabla_{h_{a}}f\nabla_{h_{a}}f^{\top}\right]$. To see its connection to the CRH, note that it can be written as
@@
Figure 8: After the training of a six-layer eight-head fully connected network, all six relations (Eq. (3)-(5)) hold strongly at a large weight decay (fc2). For a small weight decay, at least one forward and backward relation holds strongly. The shaded region shows the variation across five hidden layers, and the solid lines show the median of these alignments. At a small $\gamma$, the best alignment is between $H_{a}$ and $W^{\top}W$ for the forward relation, and $G_{a}$ and $WW^{\top}$ for the backward, in agreement with the theoretical prediction.
@@ 8 Conclusion
In this work, we propose the Canonical Representation Hypothesis (CRH), a new perspective for studying the formation of representations in neural networks. The CRH suggests that representations align with the weights and gradients after training. It is a generalization of the neural collapse phenomenon for any fully connected layer in a neural network. In this view, representations are formed based on the degree and modes of deviation from the CRH. This deviation leads to the Polynomial Alignment Hypothesis (PAH), which posits that when the CRH is broken, distinct phases emerge in which the representations, gradients, and weights become polynomial functions of each other. A key future direction is to understand the conditions that lead to each phase and how these phases affect the behavior and performance of models. The CRH may also have algorithmic implications. If representations align with the gradients (as in RGA), it might be possible to manually inject noise into neuron gradients to engineer specific structures in the model’s representations. However, the CRH and PAH have several limitations. They apply only to fully connected layers, and a future step is extending them to convolutional layers. Additionally, our research has primarily focused on characterizing the final stage of representation formation. A more comprehensive theory of representation dynamics could lead to better training algorithms.
@@ Appendix A More Related Works
The neural representation is a central object of study in both AI and neuroscience (Esser et al.,, 2020; Wang et al.,, 2018). The widespread use of techniques like t-SNE (Van der Maaten and Hinton,, 2008) for analyzing the penultimate layer representation in deep learning highlights the importance of understanding representation space. However, the theory of how the representations are formed in neural networks is scarce. Liu et al., (2022) proposes an effective theory to study representation learning in modular addition. A recent work studies representation formation from the angle of symmetries (Ziyin,, 2024), showing that permutation symmetries in the latent layers lead to neuron merging and low rankness in the representation. Ziyin et al., (2024) showed that after training, the latent representation of a deep linear net is aligned with a linear transformation of the prediction residual. A recent work showed that the evolution of the representation during the initial stage of training has universal properties shared by different types of models (van Rossem and Saxe,, 2024). Empirically, Huh et al., (2024) showed that the neural representations learned by different networks are essentially similar. Also, an emergent field in neuroscience studies how the representations learned by neural networks closely resemble those of biological neurons (Rajalingham et al.,, 2018). These results suggest the existence of some fundamentally shared mechanisms of the formation of representations. Also, neural collapse has been found to improve the transfer learning of neural networks Galanti et al., (2021), and an interesting future step may be understanding how CRH might improve transfer learning.
@@ B.2 Proof of Theorem 2
We need a few lemmas.
@@ B.3 Proof of Theorem 1
Assume assumption 1. If $\mathbb{E}[\Delta(h_{a}h_{a}^{\top})]=0$ and $\mathbb{E}[\Delta(h_{b}h_{b}^{\top})]=0$, then,
@@
Assume assumption 1. If $\mathbb{E}[\Delta(g_{a}g_{a}^{\top})]=0$ and $\mathbb{E}[\Delta(g_{b}g_{b}^{\top})]=0$, then,
@@ B.4 Proof of Proposition 1
$0=\ell(f(h+\epsilon\hat{n}))-\ell(f(h))=\epsilon\hat{n}^{\top}\nabla_{h}\ell(h).$ (113)
One can derive the other two relations simply using the definition of CRH: $\mathbb{E}[hh^{\top}]\propto\mathbb{E}[gg^{\top}]\propto Z$.
@@
$\ell(f(h+\epsilon\hat{n}))-\ell(f(h))=\epsilon\hat{n}^{\top}g+O(\epsilon).$ (115)
@@
B.5 Proof of Corollary 1
Proof.
We first prove the forward direction: $f(x)$ is invariant implies $h(x)$ is invariant. Assume that $f(x)$ is invariant to a change in $x$: $x\to x+\epsilon\hat{n}$, where $\hat{n}$ is a vector and $\epsilon$ a scalar. This implies that
$f(x+\epsilon\hat{n})=f(x)+O(\epsilon^{2}).$ (116)
Now, if $h_{a}(x+\epsilon\hat{n})=h_{a}(x)+O(\epsilon^{2})$, we are done. Thus, it suffices to consider the case where
$h_{a}(x+\epsilon\hat{n})=h_{a}^{\prime}(x)+\epsilon v+O(\epsilon^{2}),$ (117)
for some vector $v\neq 0$. But because of Eq. (116), we have
$\ell(h_{a}+\epsilon v+O(\epsilon^{2}))=\ell(h_{a})+O(\epsilon^{2}).$ (118)
By the CRH and Proposition 1, this implies that
$\mathbb{E}[h_{a}h_{a}^{\top}]v\propto\mathbb{E}[W^{\top}W]v=0.$ (119)
This implies that
$Wv=0,$ (120)
and so
$h_{b}=W(h_{a}+v)=0.$ (121)
The reverse direction is trivially true. This completes the proof. ∎
B.6 Proof of Theorem 3
@@
where we have used the assumption that $G_{b}=\mathbb{E}[\nabla_{f}\ell\nabla_{f}^{\top}\ell]=0$.
@@ C.1 Experimental Details
ResNet-18 for self-supervised learning tasks with the CIFAR-10/100 datasets. The model is the same as res1, except that the last fc layer output becomes $128$-dimensional, which is known as the projection dimension in SSL. The training follows the default procedure in the original paper (Chen et al.,, 2020), proceeding with a batch size of $512$ and $\gamma=5\times 10^{-5}$ for $1000$ epochs.
@@
a six-layer transformer ($100$M parameters) trained on the OpenWebText (OWT) dataset (Gokaslan and Cohen,, 2019); the number of parameters of this model matches the smallest version of GPT2. The model has six layers with eight heads per layer, having $100$M trainable parameters in total. For each experiment, we train with Adam with a weight decay strength of $1\times 10^{-4}$ for $10^{5}$ iterations, when the training loss stops changes significantly. Since every representation has three dimensions: data $N$, token $T$, and feature $F$, we treat each token as if they are a separate sample in computing the covariances. Namely, we contract the representation tensor along the data and token dimension, resulting in a $F\times F$ covariance matrix.

### v2

@@ 7.2 Feature-Importance Alignment and Invariant Learning
Another direct interpretation of $\mathbb{E}[gg^{\top}]$ is the saliency of a latent feature. The quantity $\mathbb{E}[gg^{\top}]$ measures how the loss function changes as a neuron activation is changed by a small degree and is often used on the input layer to measure the importance of a feature (Selvaraju et al., 2017; Adadi & Berrada, 2018). This interpretation justifies many heuristics for understanding neural networks: take the principal components of a hidden layer and study its largest eigenvalues. If there is no direct link between the magnitude of the eigenvalues in these latent representations and their importance in affecting the learning (measured by the training loss), these analyses will be unreasonable.
@@
In some sense, Proposition 1 can be seen as a generalization of NC (next section) because it applies to both regression and classification tasks. Note that there are two ways for the condition $\ell(f(h+\epsilon\hat{n}))=\ell(f(h))$ to be satisfied: (1) $f(h+\epsilon\hat{n})=f(h)$, which means that when the model output itself invariant to such a variation, the latent representation in this direction will vanish; (2) $\ell(f(h)+\epsilon\hat{n}^{T}\nabla f(h))=\ell(f(h))$, which means that any variation that does not change the loss function value will vanish. This fact also means that the model will learn a compact representation: any irrelevant latent space will have zero variation. This is consistent with the often observed matching between the rank of the representation and the inherent dimension of the problem (Papyan, 2018; Ziyin, 2024). A major phenomenon of well-trained CNNs is that the latent representations learn to become essentially invariant to task-irrelevant features (Zeiler & Fergus, 2014; Selvaraju et al., 2017), an observation that matches the high-level features of the human visual cortex (Booth & Rolls, 1998). The CRH thus suggests a reason for these observed invariances.
@@ 7.3 Reduction to Neural Collapse in Perfect Classification
We now prove that NC is equivalent to the CRH in an interpolating classifier. We focus on the first two properties here because NC3-4 are essentially consequences of NC1-2 (Rangamani & Banburski-Fahey, 2022). NC1 states that the inner-class variations of the penultimate representation vanishes. NC2 states that the average representation $\mu_{c}$ of the class $c$ is orthogonal to each other: $\mu_{c}^{\top}\mu_{c^{\prime}}=\delta_{cc^{\prime}}$. The ground truth model must be invariant to inner class variations in a classification task. Let $c$ denote the index of a class, it should be the case that the ground truth model $f^{\prime}$ satisfies: $f^{\prime}_{c}(x_{c})=\zeta\mathbf{1}_{c}$, where $\mathbf{1}_{c}$ is a one hot vector on the $c$-th dimension and $\zeta>0$ is an arbitrary scalar. Proposition 1 thus suggests that any model that recovers the ground truth must have such invariance in the output and in the latent representation $h$, which is exactly NC1.
@@ 7.4 Equivariance of CRH and Neural Feature Ansatz
A closely related hypothesis is the NFA, which is related to the feature learning in fully connected layers (Radhakrishnan et al., 2023; Beaglehole et al., 2023). Using our notation, the NFA states $W^{\top}W\propto\mathbb{E}\left[\nabla_{h_{a}}f\nabla_{h_{a}}f^{\top}\right]$. To see its connection to the CRH, note that it can be written as
@@
Figure 8: After the training of a six-layer fully connected network, all six relations (Eq. (3)-(5)) hold strongly at a large weight decay (fc2). For a small weight decay, at least one forward and backward relation holds strongly. The shaded region shows the variation across five hidden layers, and the solid lines show the median of these alignments. At a small $\gamma$, the best alignment is between $H_{b}$ and $W^{\top}W$ for the forward relation, and $G_{a}$ and $WW^{\top}$ for the backward, in agreement with the theoretical prediction.
@@ 8 Conclusion
In this work, we propose the Canonical Representation Hypothesis (CRH), a new perspective for studying the formation of representations in neural networks. The CRH suggests that representations align with the weights and gradients after training. It is a generalization of the neural collapse phenomenon for any fully connected layer in a neural network. In this view, representations are formed based on the degree and modes of deviation from the CRH. This deviation leads to the Polynomial Alignment Hypothesis (PAH), which posits that when the CRH is broken, distinct phases emerge in which the representations, gradients, and weights become polynomial functions of each other. A key future direction is to understand the conditions that lead to each phase and how these phases affect the behavior and performance of models. The CRH may also have biological implications as it implies that neural networks tend to learn an orthogonalized representation, which has been observed in the biological brain recently (Sun et al., 2025). The CRH may also have algorithmic implications. If representations align with the gradients (as in RGA), it might be possible to manually inject noise into neuron gradients to engineer specific structures in the model’s representations. However, the CRH and PAH have several limitations. They apply only to fully connected layers, and a future step is extending them to other types of layers. Additionally, we have primarily focused on characterizing the final stage of representation formation. A more comprehensive theory of representation dynamics could lead to better training algorithms.
@@ Appendix A More Related Works
The neural representation is a central object of study in both AI and neuroscience (Esser et al., 2020; Wang et al., 2018). The widespread use of techniques like t-SNE (Van der Maaten & Hinton, 2008) for analyzing the penultimate layer representation in deep learning highlights the importance of understanding representation space. However, the theory of how the representations are formed in neural networks is scarce. A recent work approached this problem from the angle of symmetries (Ziyin, 2024), showing that permutation symmetries in the latent layers lead to neuron merging and low rankness in the representation. Ziyin et al. (2024b) showed that after training, the latent representation is aligned with a linear transformation of the prediction residual. A recent work showed that the evolution of the representation during the initial stage of training have universal properties shared by different types of models (van Rossem & Saxe, 2024). Empirically, Roeder et al. (2021) and Huh et al. (2024) showed that the neural representations learned by different networks are essentially similar, which may offer a partial explanation for why the CRH seems to hold across different architectures. Also, an emergent field in neuroscience studies how the representations learned by neural networks closely resemble those of biological neurons (Rajalingham et al., 2018). These results suggest the existence of some fundamentally shared mechanisms of the formation of representations.
The finding that the neural networks find invariant structures is interesting and should be contrasted with the finding that sometimes the neural networks learns an invertible function (Zimmermann et al., 2021; Reizinger et al., 2024). This may have to do with the fact that both neural collapses have been found to emerge when there is nonnegligible strength of regularization (such as weight decay (Rangamani et al., 2021)). This suggests that some form of simplicity bias is required to make neural networks prefer invariant structures, which may be another interesting topic for future research.
@@ B.2 Proof of Theorem 2
We need a few lemmas. The following lemmas are quite general and apply to arbitrary symmetric matrices $A$ and $B$, which is clear from the lemma statement. Eventually, when we apply these lemmas, the specific $A$ and $B$ will be the $H$, $G$, $Z$ matrices we defined in the CRH statement.
@@ B.3 Proof of Theorem 1
Assume assumption 1. If $\mathbb{E}[\Delta H_{a}]=0$ and $\mathbb{E}[\Delta H_{b}]=0$, then,
@@
Assume assumption 1. If $\mathbb{E}[\Delta G_{a}]=0$ and $\mathbb{E}[\Delta G_{b}]=0$, then,
@@ B.4 Proof of Proposition 1
$\ell(f(h+\epsilon\hat{n}))-\ell(f(h))=\epsilon\hat{n}^{\top}\nabla_{h}\ell(h)+O(\epsilon^{2})=O(\epsilon^{2}),$ (113)
which is possible only if $\hat{n}^{\top}\nabla_{h}\ell(h)=0$. One can derive the other two relations simply using the definition of CRH: $\mathbb{E}[hh^{\top}]\propto\mathbb{E}[gg^{\top}]\propto Z$.
@@
$\ell(f(h+\epsilon\hat{n}))-\ell(f(h))$ $=\epsilon\hat{n}^{\top}g+O(\epsilon^{2})$ (115)
$=O(\epsilon^{2}).$ (116)
@@
B.5 Proof of Theorem 3
@@
where we have used the assumption that $G_{b}=\mathbb{E}[\nabla_{f}\ell\nabla_{f}^{\top}\ell]=I$.
@@ C.1 Experimental Details
ResNet-18 for self-supervised learning tasks with the CIFAR-10/100 datasets. The model is the same as res1, except that the last fc layer output becomes $128$-dimensional, which is known as the projection dimension in SSL. The training follows the default procedure in the original paper (Chen et al., 2020), proceeding with a batch size of $512$ and $\gamma=5\times 10^{-5}$ for $1000$ epochs.
@@
a six-layer transformer ($100$M parameters) trained on the OpenWebText (OWT) dataset (Gokaslan & Cohen, 2019); the number of parameters of this model matches the smallest version of GPT2. The model has six layers with eight heads per layer, having $100$M trainable parameters in total. For each experiment, we train with Adam with a weight decay strength of $1\times 10^{-4}$ for $10^{5}$ iterations, when the training loss stops changes significantly. Since every representation has three dimensions: data $N$, token $T$, and feature $F$, we treat each token as if they are a separate sample in computing the covariances. Namely, we contract the representation tensor along the data and token dimension, resulting in a $F\times F$ covariance matrix.
@@ C.9 Stationarity of the covariance matrix
C.10 Full Figure to Figure 4
See Figure 22.
Figure 22: This is the same figure as Figure 4. The legend shows the weight decay value used for each experiment. Dots with the same color come from the five different hidden layers for the model trained at this specific weight decay value.
Appendix D Meaning of the PAH
D.1 Linear Regression and Phases of the PAH
In some sense, the phases of the PAH already appear implicitly in many standard models. Consider a linear regression problem:
$W^{*}=\min_{W}\mathbb{E}\|Wx-y\|^{2},$ (132)
where $\mathbb{E}$ denotes averaging over the training set. Its solution is well known:
$(W^{*})^{\top}=\mathbb{E}[xx^{\top}]^{-1}\mathbb{E}[xy].$ (133)
From the perspective of the PAH, this solution can be seen as a composition of two functional layers: $W^{*}=V_{1}V_{2}$, where the first layer is $V_{1}=\mathbb{E}[xx^{\top}]^{-1/2}$ and the second layer is $V_{2}=\mathbb{E}[\tilde{x}y]$, and $\tilde{x}=\mathbb{E}[xx^{\top}]^{-1/2}x$.
The first layer normalizes the input distribution and is apparently related to the PAH because one can identify
$H_{a}=\mathbb{E}[xx^{\top}]$ (134)
as the input representation to the layer, and so
$\mathbb{E}[xx^{\top}]^{-1}=(W^{*})^{\top}W^{*}=Z_{a},$ (135)
which implies that
$H_{a}^{-1}=Z_{a},$ (136)
which can be identified as the 7th phase of the PAH in Table 1.
D.2 Potential Implications of the PAH
Here, we discuss some possible and interesting meanings of the phases of the PAH. Validating these intuitions could be of great interest.
One way the phase could imply, as we suggested in the manuscript, is that a positive relation between $H_{a}$ and $Z_{a}$ could imply a low-rank mechanism, where larger eigenvalues of $H_{a}$ are expanded while smaller eigenvalues are suppressed. The fact that neural networks learn these low-rank representations is consistent with common observation. For example, see Kobayashi et al. (2024). Another implied mechanism is representation normalization. This happens when $H_{a}\propto Z_{a}^{-1}$, which means that $H_{b}$ will be normalized after this weight. See the discussion about linear regression in Section D.1 for how this phase exists implicitly, even in linear regression.
Another kind of interesting phase is where the alignment condition implies gradient vanishing or explosion problems, which have been well-known for decades (Kanai et al., 2017; Hochreiter, 1998). Noticing that $g_{a}=W^{T}g_{b}$, it is naturally the case that if $Z_{b}$ and $G_{b}$ are aligned with a positive exponent, the gradient in the previous layer will face an explosion and vanishing problem simultaneously (as larger eigenvalues become larger). In contrast, the phase $Z_{b}\propto G_{b}^{-1}$ seems to be the ideal training phase as it implies that the gradient in the previous will be normalized. Identifying the causes of these phases is also of great future interest.